OpenStudy (idealist10):

Integrate x/(x^2+y^2)^(3/2) dx.

OpenStudy (idealist10):

@aaronq @Loser66

OpenStudy (loser66):

\[\int \dfrac{x}{\sqrt{(x^2+y^2)^3}}dx\\=\int\dfrac{x}{(x^2+y^2)\sqrt{x^2+y^2}}dx\] Let \(u=\sqrt{x^2+y^2}\) \(u^2= x^2+y^2\) and du =\(\dfrac{x}{\sqrt{x^2+y^2}}dx\) Plug all in \[\int \dfrac{du}{u^2}= \int u^{-2}du\] I think you can handle it from here.